Tohban Report 2010-12-29
From RHESSI Wiki
|Start Date:||22 Dec 2010|
|End Date:||29 Dec 2010|
|Next Tohban:||Pascal Saint-Hilaire|
|List all reports|
It was a very quiet week, with only 4 B-class flares! GOES base level has risen slightly over the last week, and there is a new active region rotating on the disk now, so perhaps there will be some activity soon.
How many GOES flares occurred?
Flares above B, C, M, X class were 4 0 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 3 0 0 0
Once again the tohban had a dream of major X-class flares, but none appeared...
One 10-min gap in SOH and monitor rates on the 23rd:
GAP START TIME GAP END TIME GAP (SEC) 2010-12-23T10:30:00.000 -- 2010-12-23T10:40:00.000 600.00000
The data has been filled in but shows a dropout; perhaps a bad packet.
SSR is emptying after each pass, with a max of 22%. Cold plate 2 is at 98.8 K. A spinup was performed yesterday, so the spin rate is at 15 rpm now. The threshold on Detector 6 was raised last Wednesday due to the occurrence of noise. For this reason, Detc 6 was also removed from the set of detectors used to calculate livetime, and therefore control the attenuator movements. This set changed from det [3,4,6,9] to det [1,3,4,9].